OpenStudy (anonymous):

Rewrite the expression with positive exponents: 1/(3x-3)

OpenStudy (phi):

you should use ^ to show an exponent. and put the exponent in parens so your problem (guessing!) is \[ \frac{1}{3x^{-3}} \] or \[ \frac{1}{(3x)^{-3}} \]

OpenStudy (anonymous):

thats what ive been doing god y cant i get things :(

OpenStudy (phi):

of the two, posted above which is your problem? (first or second?)

OpenStudy (anonymous):

first

OpenStudy (anonymous):

?

OpenStudy (phi):

if it is the first, then you could write it as 1/(3x^(-3)) the first thing is write it as (1/3) * 1/x^(-3) \[ \frac{1}{3}\cdot \frac{1}{x^{-3}} \] now use the rule that 1/x^(-n) = x^n (flip and change the sign of the exponent)

OpenStudy (anonymous):

ok

OpenStudy (phi):

people came up with this great idea x*x*x*x is too hard to write, so say x^4 they decided if you want to write \[ \frac{1}{x*x*x*x} = x^{-4} \] use a minus 4 to show the x is in the bottom if we flip both sides we get \[ \frac{x*x*x*x}{1} = \frac{1}{x^{-4}} \] of course, the left side is the same as x^4

OpenStudy (phi):

\[ x^4= \frac{1}{x^{-4}} \] can you use the same "rule" for 1/x^(-3) ?